# Fourier dimension at $C^{1+\alpha}$ regularity

### v2

@@ 3. Discretised sum-product bounds under light tails
Our strategy is to reduce the Fourier decay bound to exponential sums of multiplicative random variables coming from products of the derivatives $g_{a}^{\prime}(x)$ of the IFS. To bound this, as in Bourgain and Dyatlov in [13], one can reduce this to a multiscale non-concentration bound via the discretised sum-product phenomenon [12] coming from additive combinatorics. However, to gain the multiscale non-concentration, previous works require smoothness from the underlying dynamics such as in the form of application of Dolgopyat’s method [41, 10, 9, 27, 30]. The main novelty of our paper lies in the analysis of the non-concentration bound so it can be studied also in the low regularity that cover $C^{1+\alpha}$ systems.
@@
Firstly, let us recall a quantitative version of this sum-product phenomenon, where this version is a consequence of [36, Theorem 1.14], see also [13]:
@@ 5. Reducing multiscale non-concentration to larger probability space
We have, by Cauchy–Schwarz applied to the sum on $\mathbf{a},\mathbf{b},\mathbf{c}$:
@@ 6.2. $C^{2}$ IFSs: Using UNI to prove the non-concentration
The goal of this subsection is to prove 1.2, that is, we will show that (QNL) holds under (UNI).
To this aim, we will state and prove a countable $C^{1+\alpha}$ branch version of the Tree Lemma presented in [28, Lemma 2.5.7] in the $C^{2}$ finite branch setting. This type of “tree Lemma” is useful as it allows us to avoid the use of spectral gap argument while still adapting some parts of a Dolgopyat style approach. The main difficulty of the countable compared to [28, Lemma 2.5.7] is the possible blow-ups of derivatives caused by the parabolicity or infinite number of branches, which prevents us to use same parameters as in the finite branch case. While we cannot work with the derivatives $\partial_{t}\log\left|g_{\mathbf{a}}^{\prime}(t)\right|_{|t=x}$ directly, it turns out that the simplicity of the tree Lemma allows us to conclude the necessary conclusion with the following $\alpha$-Hölder difference quotients of $\log\left|g_{\mathbf{a}}^{\prime}\right|$:
@@
To prove Proposition 6.6, we need first the following important non-concentration on the fractal $F$ where UNI is used.
@@
Assume the hypotheses of Proposition 6.6. Suppose $0<p<\frac{c_{0}}{4}$ is a fixed parameter and $t:F^{2}\to\mathbb{R}$. Then for every $c\in\mathcal{A}$ the following holds true: If $\mathbf{a}\in\Sigma^{N}$ and $\mathbf{b}\in\Sigma^{n-N}$ with $b\left(\mathbf{a}\right)=b\left(\mathbf{b}\right)=c,\mathbf{b}^{\prime}\mathbf{a}\in\Sigma^{n}$, then
$\forall x,y\in F\cap I_{c}:x\neq y\implies\lambda_{\mathbf{b}^{\prime}\mathbf{a}}^{\alpha}\left(x,y\right)\not\in\left[t\left(x,y\right)-p,t\left(x,y\right)+p\right]$
@@
Let $x,y\in F\cap I_{c}$ be fixed and $t:F^{2}\to\mathbb{R}$ a map. One of the words, call it $\mathbf{c}$, satisfies
$\left|\lambda_{\mathbf{c}}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|\geq\frac{c_{0}}{2}.$
Otherwise we would arrive at a contraction in (6.2) after adding and subtracting $t\left(x,y\right)$ and using the triangle inequality. We are ultimately looking to show that the inequality
@@
holds for any prefix $\mathbf{b}\in\Sigma^{n-N}$. Let $\mathbf{b}\in\Sigma^{n-N}$ be fixed, but arbitrary. By the chain rule we obtain that
@@
7. Proof of the Measure Non-Linearity condition for $(\Phi_{\mu},-\delta_{\mu}\tau_{\mu})$
This section is devoted to the proof of Theorem 1.4. Recall some of the notations: we have fixed $C\geq 1$ and $\varepsilon>0$ and denoted
$\mathcal{M}_{C,\varepsilon}=\{\mu\in\mathcal{P}([-1,1]):\forall\text{ interval }I\text{ with }|I|\leq\tfrac{1}{100}\text{ we have }\mu(I)\leq C|I|^{\varepsilon}\}$
Given $\mu\in\mathcal{M}_{C,\varepsilon}$, define the branches of an IFS $\Phi_{\mu}$ by
$g_{\mu,1}(x):=\kappa\int_{0}^{x}\exp(\mu(-\infty,t])\,dt+b_{1},\quad x\in[0,1],$
and $g_{\mu,0}(x)=\kappa_{0}x+b_{0}$, where $0<\kappa_{0}\leq\kappa<1/10$, $b_{0},b_{1}\in\mathbb{R}$ is chosen such that the IFS $\Phi_{\mu}=\{g_{0,\mu},g_{1,\mu}\}$ satisfies the strong separation condition. Let $T_{\mu}$ be the associated expanding map on the attractor $F_{\mu}$ to $\Phi_{\mu}$, and write the geometric potential $\tau_{\mu}(x):=-\log|T_{\mu}^{\prime}(x)|$. Note by the Definition of each $\mu\in\mathcal{M}_{C,\varepsilon}$, the IFS $\Phi_{\mu}$ is $C^{1}$ and $\tau_{\mu}$ is $\alpha$-Hölder map for any $\alpha\leq\varepsilon$. Finally, we denoted the $\mu$-controlled transfer operators
$\mathcal{L}_{\mu}f(x):=\sum_{i}|g_{\mu,i}^{\prime}(x)|^{\delta_{\mu}}f(g_{\mu,i}(x)),$
@@ 7. Proof of the Measure Non-Linearity condition for $(\Phi_{\mu},-\delta_{\mu}\tau_{\mu})$
where $\delta_{\mu}\in(0,1)$ is chosen as the unique number such that

### v3

@@ 3. Discretised sum-product bounds under light tails
Our strategy is to reduce the Fourier decay bound to exponential sums of multiplicative random variables coming from products of the derivatives $g_{a}^{\prime}(x)$ of the IFS. To bound this, as in Bourgain and Dyatlov in [18], one can reduce this to a multiscale non-concentration bound via the discretised sum-product phenomenon [17] coming from additive combinatorics. However, to gain the multiscale non-concentration, previous works require smoothness from the underlying dynamics such as in the form of spectral gaps of twisted transfer operators [67, 13, 11, 43, 47], which we can avoid.
@@
Firstly, let us recall a quantitative version of this sum-product phenomenon, where this version is a consequence of [60, Theorem 1.14], see also [18]:
@@ 5. Reducing multiscale non-concentration to larger probability space
We have, by Cauchy-Schwarz applied to the sum on $\mathbf{a},\mathbf{b},\mathbf{c}$:
@@ 6.2. $C^{2}$ IFSs: Using UNI to prove the non-concentration
The goal of this subsection is to prove UNI part of 1.2, that is, we will show that (QNL) holds under (UNI).
To this aim, we will state and prove a countable $C^{1+\alpha}$ branch version of the Tree Lemma presented in [44, Lemma 2.5.7] in the $C^{2}$ finite branch setting. This type of “tree Lemma” is useful as it allows us to avoid the use of spectral gap argument while still adapting some parts of a Dolgopyat style approach. The main difficulty of the countable compared to [44, Lemma 2.5.7] is the possible blow-ups of derivatives caused by the parabolicity or infinite number of branches, which prevents us to use same parameters as in the finite branch case. While we cannot work with the derivatives $\partial_{t}\log\left|g_{\mathbf{a}}^{\prime}(t)\right|_{|t=x}$ directly, it turns out that the simplicity of the tree Lemma allows us to conclude the necessary conclusion with the following $\alpha$-Hölder difference quotients of $\log\left|g_{\mathbf{a}}^{\prime}\right|$:
@@
To prove Proposition 6.6, we need first the following important non-concentration on the fractal $F$ where UNI is used. Before we can state that result, let us first review the exact $\left(c_{0},\alpha\right)$-UNI condition that we are interested in, and then introduce two useful definitions: The aforementioned $\left(c_{0},\alpha\right)$-UNI condition was:
  • (UNI)
    $|X_{\mathbf{a}}(x,y)-X_{\mathbf{b}}(x,y)|\ \geq\ c_{0}\,|x-y|^{\alpha},$
where
$X_{\mathbf{a}}(x,y):=S_{n}\tau\circ g_{\mathbf{a}}(x)-S_{n}\tau\circ g_{\mathbf{a}}(y)=\log g_{\mathbf{a}}^{\prime}(x)-\log g_{\mathbf{a}}^{\prime}(y),\qquad\mathbf{a}\to x,y.$
Next, let $c\in\mathcal{A},N\in\mathbb{N},\mathbf{a},\mathbf{b}\in\Sigma^{N}$ and $c_{0}>0,\alpha>0$ be such that $b\left(\mathbf{a}\right)=b\left(\mathbf{b}\right)=c$ and $\forall x\neq y\in F\cap I_{c}:\left|X_{\mathbf{a}}\left(x,y\right)-X_{\mathbf{b}}\left(x,y\right)\right|\geq c_{0}\left|x-y\right|^{\alpha}$.
Define such a pair $\left(\mathbf{a},\mathbf{b}\right)$ as $\left(c_{0},\alpha\right)$-UNI pair of length $N$ at leaf $c$. Next, given $N\in\mathbb{N},c\in\mathcal{A}$, define the set
[…]
@@
Assume the hypotheses of Proposition 6.6, suppose $0<p<\frac{c_{0}}{4}$ is a fixed parameter, let $t:F^{2}\to\mathbb{R}$ and assume that $\left(N_{k}\right)_{k\in\mathbb{N}}\subset\mathbb{N}$ is a sequence of integers satisfying the aforementioned $\left(c_{0},\alpha\right)$-Uniform Non-Integrability condition. Let $k\in\mathbb{N}$ be fixed, but arbitrary. Then,
$\forall k\in\mathbb{N}:\forall c\in\mathcal{A}:\exists\mathfrak{p}\in\mathrm{UP}\left(N_{k},c\right):\forall x,y\in F\cap I_{c}:x\neq y\implies$
$\bigg(\exists\mathbf{c}\left(x,y\right)\in\mathfrak{p}:\forall n\geq N_{k}:\forall\mathbf{b}\in\Sigma^{n-N_{k}}:\left[\left(b\left(\mathbf{b}\right)=c\right)\land\left(\mathbf{b}^{\prime}\mathbf{c}\left(x,y\right)\in\Sigma^{n}\right)\right]\implies$
$\lambda_{\mathbf{b}^{\prime}\mathbf{c}\left(x,y\right)}^{\alpha}\left(x,y\right)\not\in\left[t\left(x,y\right)-p,t\left(x,y\right)+p\right]\bigg)$
@@
Let $x,y\in F\cap I_{c}$ be fixed and $t:F^{2}\to\mathbb{R}$ a map. One of the words, call it $\mathbf{c}\left(x,y\right)$, satisfies
$\left|\lambda_{\mathbf{c}\left(x,y\right)}^{\alpha}\left(x,y\right)-t\left(x,y\right)\right|\geq\frac{c_{0}}{2},$
at $\left(x,y\right)$. Otherwise we would arrive at a contraction in (6.2) after adding and subtracting $t\left(x,y\right)$ and using the triangle inequality. For the rest of the proof this pair $\left(x,y\right)$ is fixed and so we write $\mathbf{c}:=\mathbf{c}\left(x,y\right)$.
We are ultimately looking to show that the inequality
@@
holds for any prefix $\mathbf{b}\in\Sigma^{n-N}$ at $\left(x,y\right)$. Let $\mathbf{b}\in\Sigma^{n-N}$ be fixed, but arbitrary. By the chain rule we obtain that
@@
7. Proof of Theorem 1.3 on the Lyons measure
In this section we will prove Theorem 1.3 by inducing approach and using Theorem 1.1. Fix $t\in(0,1)$. The Lyons IFS is the parabolic Möbius system
$\Phi=\Big\{f_{t}(x)=\frac{x+t}{x+t+1},\qquad f_{0}(x)=\frac{x}{x+1}\Big\},\qquad x\in[0,1].$
With equal weights $(1/2,1/2)$, its invariant measure $\nu_{t}$ is Lyons conductance measure with the law of a random continued fraction with i.i.d. digits in $\{0,t\}$, see [53, 72, 57]. As in [57], we accelerate along $\phi_{2}$ and obtain the induced uniformly expanding countable IFS
$\tilde{\Phi}=\{\psi_{n}^{t}=f_{0}^{n}\circ f_{t}:\ n\in\mathbb{N}_{0}\},\qquad\psi_{n}^{t}(x)=\frac{x+t}{(n+1)(x+t)+1}.$
Its branches satisfy
$(\psi_{n}^{t})^{\prime}(x)=\frac{1}{\big((n+1)(x+t)+1\big)^{2}}<1,\qquad x\in[0,1],$
and the invariance relation expands as
$\nu_{t}=\frac{1}{2}\,(\phi_{1}^{t})_{*}\nu_{t}+\frac{1}{2}\,(\phi_{2})_{*}\nu_{t}\quad\Longrightarrow\quad\nu_{t}=\frac{1}{2}\sum_{n=0}^{\infty}2^{-n}(\psi_{n}^{t})_{*}\nu_{t}.$
Thus $\nu_{t}$ is a projection of a Gibbs measure on $\mathbb{N}^{\mathbb{N}}$ associated to a locally constant potential $\phi(\mathbf{a}):=\log 2^{-a_{1}}$ on $\mathbb{N}^{\mathbb{N}}$. Thus by Lemma 2.3 $\nu_{t}$ has no atoms so we are in realm where Theorem 1.1 can be applied if we verify the UNI condition (so QNL holds by Theorem 1.2) and light tail.
Let us now verify the UNI condition for the induced IFS.
Lemma 7.1.
For every $t\in(0,1)$ the accelerated Lyons IFS $\tilde{\Phi}=\{\psi_{n}^{t}\}_{n\geq 0}$ satisfies the $(c_{0},1)$-UNI condition.
Proof.
Consider the branches
$\psi_{0}^{t}(x)=\frac{x+t}{x+t+1},\qquad\psi_{1}^{t}(x)=\frac{x+t}{2x+2t+1},$
giving us a finite uniformly hyperbolic analytic sub-IFS $\Phi_{0}:=\{\psi_{0}^{t},\psi_{1}^{1}\}$. Now, we will just verify the $(c_{0},1)$-UNI for $\Phi_{0}$, which implies it for $\tilde{\Phi}$ since if for $N\in\mathbb{N}$ we have the UNI condition for words $\mathbf{a},\mathbf{b}\in\{0,1\}^{N}$, these will also appear coding in the infinite IFS.
To verify $(c_{0},1)$-UNI, since $\Phi_{0}$ is analytic and hyperbolic, we can rely on [7, Claim 2.1] as an avenue to prove this. We say that $\Phi_{0}$ is not $C^{2}$-conjugate to a linear IFS if there does not exist a $C^{2}$ diffeomorphism $h:[0,1]\to h([0,1])\subset\mathbb{R}$ such that for all $i\in\mathcal{\{}0,1\}$,
$h\circ\varphi_{i}\circ h^{-1}(x)=a_{i}x+b_{i}$
for some constants $a_{i}\neq 0$, $b_{i}\in\mathbb{R}$. Then Claim 2.1 of [7] says if $\Phi_{0}$ is not $C^{2}$-conjugate to a linear IFS, then there exists $N\in\mathbb{N}$ and admissible words $\mathbf{a},\mathbf{b}\in\{0,1\}^{N}$ such that the corresponding branches $\psi_{\mathbf{a}}^{t},\psi_{\mathbf{b}}^{t}$ satisfy
$m\leq\left|\partial_{x}\Big(\log(\psi_{\mathbf{a}}^{t})^{\prime}(x)-\log(\psi_{\mathbf{b}}^{t})^{\prime}(x)\Big)\right|\leq m^{\prime},\qquad x\in[0,1],$
for some constants $m,m^{\prime}>0$ and $m-2\tilde{C}\,\tilde{\rho}^{\,N}>0$. This implies the UNI since for any $\mathbf{c}\in\{0,1\}^{n}$ the same estimate holds for the concatenations $\mathbf{c}\mathbf{a}$ and $\mathbf{c}\mathbf{b}$ up to a bounded distortion error:
$\left|\partial_{x}\Big(\log(\psi_{\mathbf{c}\mathbf{a}}^{t})^{\prime}-\log(\psi_{\mathbf{c}\mathbf{b}}^{t})^{\prime}\Big)(x)\right|\geq m-2\tilde{C}\,\tilde{\rho}^{\,N}.$
Indeed, using the chain rule we have
$\log(\psi_{\mathbf{c}\mathbf{a}}^{t})^{\prime}(x)=\log(\psi_{\mathbf{a}}^{t})^{\prime}(\psi_{\mathbf{c}}^{t}(x))+\log(\psi_{\mathbf{c}}^{t})^{\prime}(x),$
and similarly for $\mathbf{c}\mathbf{b}$. Subtracting gives
$\log(\psi_{\mathbf{c}\mathbf{a}}^{t})^{\prime}(x)-\log(\psi_{\mathbf{c}\mathbf{b}}^{t})^{\prime}(x)=\Big(\log(\psi_{\mathbf{a}}^{t})^{\prime}-\log(\psi_{\mathbf{b}}^{t})^{\prime}\Big)(\psi_{\mathbf{c}}^{t}(x)).$
Differentiating and using bounded distortion of $\tilde{\Phi}$ yields
$\Big|\partial_{x}\Big(\log(\psi_{\mathbf{c}\mathbf{a}}^{t})^{\prime}-\log(\psi_{\mathbf{c}\mathbf{b}}^{t})^{\prime}\Big)(x)\Big|\geq m-2\tilde{C}\tilde{\rho}^{\,N}=:c_{0}>0.$
Thus by mean value theorem then gives for all $x\neq y\in[0,1]$,
$\big|X_{\mathbf{c}\mathbf{a}}(x,y)-X_{\mathbf{c}\mathbf{b}}(x,y)\big|\geq c_{0}|x-y|,$
which is precisely the $(c_{0},\alpha)$-UNI estimate (with $\alpha=1$) required in Theorem 1.1.
Thus we just need to verify $\Phi_{0}$ is not $C^{2}$-conjugate to a linear IFS. This will follow if we prove the following non-$C^{1}$-cohomology to constant property for the IFS $\Phi_{0}$: there does not exist $u\in C^{1}$ and constants $c_{0},c_{1}$ with
$\log|(\psi_{j}^{t})^{\prime}(x)|=u(\psi_{j}^{t}(x))-u(x)+c_{j},\quad j=0,1.$ (7.1)
Indeed, arguing by contraposition, suppose that $\Phi_{0}$ were $C^{2}$-conjugate to a linear IFS. Then there exists a $C^{2}$ diffeomorphism $h:[0,1]\to h([0,1])\subset\mathbb{R}$ and constants $a_{j}\neq 0$, $b_{j}\in\mathbb{R}$ such that
$h\circ\psi_{j}^{t}\circ h^{-1}(y)=a_{j}y+b_{j},\qquad j=0,1.$
Equivalently, $\psi_{j}^{t}=h^{-1}\circ L_{j}\circ h$ where $L_{j}(y)=a_{j}y+b_{j}$. Differentiating and using the chain rule gives for all $x\in[0,1]$
$(\psi_{j}^{t})^{\prime}(x)=a_{j}\,\frac{h^{\prime}(x)}{h^{\prime}(\psi_{j}^{t}(x))}.$
Taking logarithms and setting $u(x):=-\log h^{\prime}(x)\in C^{1}([0,1])$ and $c_{j}:=\log|a_{j}|$, we obtain
$\log|(\psi_{j}^{t})^{\prime}(x)|=u(\psi_{j}^{t}(x))-u(x)+c_{j},\qquad j=0,1,$
which is precisely (7.1). Therefore the failure of (7.1) implies that $\Phi_{0}$ cannot be $C^{2}$-conjugate to a linear IFS.
So we are done if we can prove the non-$C^{1}$-cohomology to constant property (7.1) for $\Phi_{0}$. Assume on the contrary that (7.1) holds. Consider the composed branch
$\psi_{(1,0)}^{t}(x)=\frac{(1+t)x+t^{2}+2t}{(3+2t)x+2t^{2}+5t+1}.$
Let $x_{0},x_{1}$ be the positive fixed points of $\psi_{0}^{t},\psi_{1}^{t}$ and let $x_{10}$ be the fixed point of $\psi_{(1,0)}^{t}$:
$x_{0}=\frac{-t+\sqrt{t^{2}+4t}}{2},\qquad x_{1}=\frac{-t+\sqrt{t^{2}+2t}}{2},\qquad x_{10}=\frac{-t(1+t)+\sqrt{t(1+t)(t+2)(t+3)}}{3+2t}.$
Then by (7.1) we know that
$\log(\psi_{(1,0)}^{t})^{\prime}(x_{10})=\log(\psi_{0}^{t})^{\prime}(x_{0})+\log(\psi_{1}^{t})^{\prime}(x_{1}).$
Indeed, summing (7.1) along the periodic orbit of the word $(1,0)$ (at $x_{10}$) yields
$\log(\psi_{(1,0)}^{t})^{\prime}(x_{10})=c_{1}+c_{0}.$
Evaluating (7.1) at $x_{0},x_{1}$ gives $\log(\psi_{0}^{t})^{\prime}(x_{0})=c_{0}$ and $\log(\psi_{1}^{t})^{\prime}(x_{1})=c_{1}$ as claimed.
However, if we define
$Q(t):=\log(\psi_{(1,0)}^{t})^{\prime}(x_{10})-\log(\psi_{0}^{t})^{\prime}(x_{0})-\log(\psi_{1}^{t})^{\prime}(x_{1})$
we can compute
$(\psi_{0}^{t})^{\prime}(x)=\frac{1}{(x+t+1)^{2}},\qquad(\psi_{1}^{t})^{\prime}(x)=\frac{1}{(2x+2t+1)^{2}},$
$(\psi_{(1,0)}^{t})^{\prime}(x)=\frac{1}{\big((3+2t)x+2t^{2}+5t+1\big)^{2}}.$
so evaluating the derivatives at these periodic points yields
$(\psi_{0}^{t})^{\prime}(x_{0})$ $=\frac{4}{\big(t+2+\sqrt{t^{2}+4t}\big)^{2}},$
$(\psi_{1}^{t})^{\prime}(x_{1})$ $=\frac{1}{\big(t+1+\sqrt{t^{2}+2t}\big)^{2}},$
$(\psi_{(1,0)}^{t})^{\prime}(x_{10})$ $=\frac{1}{\big(t^{2}+3t+1+\sqrt{t(1+t)(t+2)(t+3)}\big)^{2}}.$
Substituting gives
$Q(t)=2\log\!\Bigg(\frac{\big(t+1+\sqrt{t^{2}+2t}\big)\big(t+2+\sqrt{t^{2}+4t}\big)}{2\big(t^{2}+3t+1+\sqrt{t(1+t)(t+2)(t+3)}\big)}\Bigg),$ (7.2)
but $Q(t)\neq 0$ for all $t\in(0,1)$. This is a contradiction. Thus (7.1) holds. ∎
To finish the proof of Theorem 1.3, we will just need to check that satisfies the light tail property of $\nu_{t}$.
Lemma 7.2.
The conductance measure $\nu_{t}$ is light-tailed for the accelerated IFS for all $\gamma\in\mathbb{R}$, that is, for where $p_{n}=2^{-(n+1)}$, $n\geq 0$: $\forall\gamma\in\mathbb{R}:\quad\sup_{x\in[0,1]}\sum_{n=0}^{\infty}p_{n}\,\big|(\psi_{n}^{t})^{\prime}(x)\big|^{-\gamma}<\infty.$
Proof.
For the light tail, note that
$(\psi_{n}^{t})^{\prime}(x)=\frac{1}{\big((n+1)(x+t)+1\big)^{2}},$
so for any $x\in[0,1]$
$\sum_{n=0}^{\infty}p_{n}\,\big|(\psi_{n}^{t})^{\prime}(x)\big|^{-\gamma}=\sum_{n=0}^{\infty}\frac{1}{2^{n+1}}\big((n+1)(x+t)+1\big)^{2\gamma}\leq\sum_{n=0}^{\infty}\frac{1}{2^{n+1}}\big((n+1)t+1\big)^{2\gamma}=:S.$
If $\gamma\leq 0$, then $S\leq 1$. If $\gamma>0$, the ratio test yields $S<1$ as
$\frac{(n+2)^{2\gamma}/2^{n+2}}{(n+1)^{2\gamma}/2^{n+1}}=\frac{1}{2}\Big(\frac{n+2}{n+1}\Big)^{2\gamma}\longrightarrow\frac{1}{2}<1.$
Thus Theorem 1.3 is proved by Theorem 1.1.
8. Proof of Theorem 1.4 on Patterson-Sullivan measures
Theorem 1.4 on Patterson-Sullivan measures for hyperbolic surfaces with cusps follows from Theorem 1.1 when coding the limit set with the coding by Li and Pan [48, Prop. 4.1]. Indeed Li-Pan construct a symbolic coding in which $\mu_{\Gamma}$ is a Gibbs measure with a light-tail property for a subshift of finite type of an IFS $\Phi$ satisfying (1)–(3) in our setting, and $\Lambda_{\Gamma}$ coincides with the attractor $F_{\Phi}$. More precisely, there exists a fundamental interval $I\subset\dot{\mathbb{R}}$ for the parabolic fixed point at $\infty$ (denoted $\Delta_{\infty}$ in [48, §2.3]), constants $C_{1}>0$, $\kappa_{+}\in(0,1)$, $\gamma_{0}>0$, a countable family of disjoint open sets $\{I_{a}\}_{a\in\mathcal{A}}\subset I$, and an expanding map $T:\bigcup_{a}I_{a}\to I$ such that:
  • (1)
    $\sum_{a\in\mathcal{A}}\mu_{\Gamma}(I_{a})=\mu_{\Gamma}(I)$;
  • (2)
  • (3)
    (uniform contraction) $|g_{a}^{\prime}(x)|\leq\kappa_{+}$ for $x\in I$;
  • (4)
  • (5)
With $\phi(x):=-\delta\log|\gamma_{a}^{\prime}(x)|_{\mathbb{S}^{1}}$ for $x\in I_{a}$, equivariance yields that $\mu_{\Gamma}$ is a Gibbs measure for $\phi$, and the $(C_{\phi,\gamma_{0}},\gamma_{0})$–tail condition holds. Finally, Li–Pan prove a UNI statement [48, Lemma 4.5]: writing $\Phi^{n}=\{\gamma_{\mathbf{a}}:\mathbf{a}\in\mathcal{A}^{n}\}$ and $\tau_{n}(x)=\sum_{k=0}^{n-1}\tau(T^{k}x)$, there exist $r,c_{0}>0$ such that for all large $N$ there are words $\{\gamma_{mj}\}_{m=1,2;\,1\leq j\leq j_{0}}$ with
which implies $(c_{0},1)$–UNI by the mean value theorem as we argued in the proof of Theorem 1.3. Therefore Theorem 1.1 applies, proving Theorem 1.4. The decay exponent in Theorem 1.4 can be traced explicitly in terms of the contraction/distortion data, the UNI constants, and the tail exponent (all available from [48]).
9. Proof of Theorem 1.5 on Manneville-Pommeau and Lorenz systems
This section is devoted to the proof of Theorem 1.5, which gives Fourier decay for some invariant measures for the non-accelerated dynamics of the Manneville-Pommeau map $T_{\alpha}$. The proof strategy is similar to Theorem 1.3 but as the measures we consider are not just Bernoulli measures like for the Lyons conductance measures $\nu_{t}$, we need more careful use of the unducing approach. The strategy here may be of independent interest to get power Fourier decay for Gibbs measures using inducing in general.
$\left|(S_{n}\tau\circ g_{\mathbf{a}})^{\prime}\right|\leq C,$ (9.1)
[…]
Lemma 9.1 (UNI).
The $T_{A}$ associated to the MP map satisfies $(c_{0},1)$-UNI on $I$.
Proof of Lemma 9.1.
$\mathbf{a}=0^{N-1}1\quad\text{and}\quad\mathbf{b}=0^{N},$
and thus
which is exactly the bound of $(c_{0},1)$-UNI. ∎
Remark.
As for UNI condition for Lorenz maps, we can argue as in the proof of Theorem 1.3. As with Theorem 1.3, the sub-IFS $\{g_{0},g_{1}\}$ for the induced system satisfies UNI if it is not $C^{2}$ conjugated to a linear IFS. Now $C^{2}$ conjugacy to an affine IFS is not possible in general, and we do this for specific numbers here to get examples of Lorenz maps. Indeed, denoting by $x_{0},x_{1},x_{10}$ the fixed points of $g_{0},g_{1},g_{10}$ respectively, we have that $\log(g_{1}\circ g_{0})^{\prime}(x_{10})=\log g_{0}^{\prime}(x_{0})+\log g_{1}^{\prime}(x_{1})$. We omit the details here, once can proceed as in Theorem 1.3 by finding the fixed points and computing the derivatives. One can see this already in simulations e.g. for $b_{0}=-1/2$, $b_{1}=1/2$, $a=1.1$ and $\alpha=0.25$ reveals $g_{0}(x)=-0.68301(0.5-x)^{4}$ with fixed point $x_{0}=-0.07429$, $g_{0}’(x)=2.73205(0.5-x)^{3}$ with $g_{0}’(x_{0})=0.51748$, $g_{1}(x)=0.68301(0.5+x)^{4}$ with fixed point $x_{1}=0.07429$, $g_{1}’(x)=2.73205(0.5+x)^{3}$ with $g_{1}’(x_{1})=0.51748$ so the product $g_{0}’(x_{0})g_{1}’(x_{1})\approx 0.26779$. Moreover, the composition $\left(g_{1}\circ g_{0}\right)(x)=0.683013(0.5-0.683013(0.5-x)^{4})^{4}$ has fixed point $x_{10}=0.03259$ and the derivative $(g_{1}\circ g_{0})’(x)=7.46411(0.5-0.683013(-0.5+x)^{4})^{3}(0.5-x)^{3}$ with $(g_{1}\circ g_{0})’(x_{10})\approx 0.077825$. Thus: $|\log(g_{1}\circ g_{0})’(x_{10})-\log g_{0}’(x_{0})-\log g_{1}’(x_{1})|\approx 1.23572$.
Thus we are in the setting of Theorem 1.1 to have that all equilibrium states for $T_{A}$ of light-tail have powefr Fourier decay. Then to go back to the parabolic system, we will just have to use the inducing structure in the following:
Proof of Theorem 1.5 assuming Theorem 1.1.
Let us first write the Manneville-Pommeau map case. Firstly, since
we can write
Then we arrive to the formula
$\|\chi_{a,k}\|_{C^{\alpha}}$ $\lesssim 1,$
$\|\log\psi_{a,k}\|_{C^{1+\alpha}}$ $\lesssim 1.$
As $\inf|\psi^{\prime}|>c$ and $|T_{\alpha}^{\prime}|\geq 1$, we have
which concludes the proof.
10. Proof of Theorem 1.6 on examples of $C^{1+\alpha}$ IFSs
We now prove Theorem 1.6 by constructing $C^{1+\alpha}$ IFSs satisfying MNL, yielding fractal equilibrium states with power Fourier decay. The construction proceeds via fixed point theory in the moduli space of IFSs parametrized by equilibrium states, ensuring the required non-concentration to implement the proof of Theorem 1.1. The key input is the existence of a branch whose log-derivative has $\alpha$–Hölder lower regularity, as furnished by a Cantor-type function with jumps precisely on the attractor.
Theorem 10.1.
@@
which depend on the equilibrium state $\mu$, where $\delta_{\mu}\in(0,1)$ is chosen as the unique number such that